# references.bib

### v2

@@ 2.3 Prediction-Powered Inference
$SE_{j,\text{Olofsson}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
@@
$C_{j,\text{Olofsson}}=(\hat{\theta}_{j,\text{Olofsson}}-1.96\cdot SE_{j,\text{Olofsson}},\ \hat{\theta}_{j,\text{Olofsson}}+1.96\cdot SE_{j,\text{Olofsson}}).$
Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$. Thus, we can also perform area estimation using methods such as PPI and stratification. If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the Olofsson confidence interval is identical to the stratified confidence interval.
@@ 3 Area estimation
In this section, we demonstrate uncertainty quantification for two area estimation use cases. First, we simulate estimating maize area in Iowa in 2022, artificially degrading the map product quality relative to simulated “ground truth” points, in order to observe the effect of map product noise and bias on the area estimate under different methods. We will see through this simulation the importance of bias correction before using maps to estimate total area. We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015. We compare the classical method, naive imputation, stratification, PPI, stratified PPI, and the method of Olofsson et al. (2013) [olofsson2013making].
@@ 3.1 Simulation: Maize area in Iowa
In each trial, we sample $n=100$ “ground truth” points uniformly at random from the $N=100000$ map product points. Initially, we set the map product and ground truth equal to each other at the 100 ground truth locations. We then experiment with degrading the map product quality as follows. We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across naive imputation, classical estimation, and PPI. We average results over 100 trials.
@@ Results
As we add increasing amounts of random noise to the map, all confidence intervals for maize area remain centered at the true maize fraction of 0.35 (Figure 5). This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same. The classical confidence interval stays constant because it only uses ground truth points. The map-based imputation confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product. However, the imputation confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
@@
When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals. This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error. As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the classical confidence interval at noise probability $p=1$. (At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth samples.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$). Thus, the effectiveness of PPI is dependent on the quality of the map product. The Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI.
@@
Figure 5: Simulation estimating maize fraction under different levels of map noise using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI results are overlapping. Left: Maize fraction 95% confidence intervals at different noise levels. Right: Maize fraction effective sample size at different noise levels. (At noise level $p$, about $p$ fraction of the points are resampled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 (the true maize fraction) and non-maize with probability 0.65.)
@@
Figure 6: Simulation estimating maize fraction 95% confidence intervals under different levels of map bias using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI confidence intervals are overlapping. (Noise is fixed at 0.4 so about 40% of the points are resampled. At bias $p$, each resampled point is labeled as maize with probability $p$ and non-maize with probability $1-p$.)
@@
The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 6. Again, the classical confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points. As bias increases, the fraction of map product points labeled as “maize” increases. As a result, the map-based imputation estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35. The imputation confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product. In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product. The PPI confidence intervals are smaller than the classical intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level). Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points. Again, the Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
@@
We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7). As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and Olofsson confidence intervals. Thus, the left panel of Figure 7 resembles an inverted version of the left panel of Figure 5.
@@
Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI confidence intervals are overlapping.
@@
As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly. (To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-based imputation estimates of maize fraction decrease linearly as model accuracy increases. In the right panel of Figure 7, a model accuracy of 0.86 corresponds to an imputed maize fraction of 0.21, significantly lower than the true fraction of 0.35. This shows that even a map with high model accuracy can yield biased estimates under naive imputation.
@@
Figure 9: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points. The Olofsson confusion matrix method, PPI, and stratification all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015. The Olofsson confusion matrix method, PPI, and stratification all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9). The Olofsson and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples. The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$. (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth. In particular, stratification is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
In this case, the naive imputation estimate — using only the map to estimate deforestation without any bias correction via ground truth labels — falls within the other confidence intervals, suggesting that the NASA Global Forest Cover Change map has small bias for estimating deforestation area. However, we emphasize that this would not have been clear without the ground truth data, and ground truth data is still necessary to check the reliability of the map product for deforestation analysis.
@@ 4 Estimating regression coefficients
In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021. We compare the classical method, naive imputation, PPI, stratification, and stratified PPI. Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
@@ Results
The general degree of uncertainty in estimating the coefficients from our non-imputation methods are as follows. The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope. Thus, deforestation is negatively associated with all covariates except for elevation. The elevation coefficient confidence intervals intersect with zero for all non-imputation methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate. The most significant improvement is for the slope covariate. Unlike in our previous area estimation examples, stratification (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers. Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads. The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
For slope and distance from rivers, the imputed coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points. The imputed coefficient for slope is about 0, and the imputed coefficient for distance from rivers is about $-0.010$. In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero. However, if we had used only the remote sensing map, we would not have discovered the negative relationship between slope and deforestation. Thus, using only the map product could result in misleading conclusions due to biases in the deforestation map product that are correlated with covariates (e.g., the map product overestimates deforestation at locations with high slope, or underestimates deforestation at locations with low slope).
Figure 11: Brazilian Amazon deforestation logistic regression coefficient 95% confidence intervals using $N=963548$ map product points and $n=1386$ ground truth points. The PPI methods result in the smallest confidence intervals, while stratification is not generally effective. For slope and distance from rivers, the map-imputed coefficient falls outside the confidence intervals for all methods that use ground truth points.
@@
Figure 13: US tree cover linear regression coefficient 95% confidence intervals using $N=983238$ map product points and $n=983$ ground truth points. PPI outperforms the classical method. For elevation, the map-imputed coefficient falls outside the PPI confidence interval.
@@
For elevation, the imputed coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval. This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations). Similar to the deforestation logistic regression use case, using only the map product could result in misleading conclusions. If we had used only the remotely sensed tree cover predictions, we would have overestimated the magnitude of the relationship between elevation and tree cover due to map errors.
@@ 5 Discussion
Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures. In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products. We find that stratified sampling, prediction-powered inference, and (for area estimation) the Olofsson method result in lower uncertainty than classical estimates using only ground truth data sampled. Unlike naive imputation using the map product alone, these estimates are guaranteed to be statistically valid.
Bias is the main issue for naive imputation.
Our experiments illustrated that a naive imputation of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias. In our simulation estimating maize fraction, naive imputation deviated from the true maize fraction of 0.35 as the bias of the crop type map increased. Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), naive imputation estimated maize fraction at 0.21 (a 40% underestimate).
In our experiments on real data, imputation for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product. However, imputation of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods. Naive imputation would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does. Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients. Similarly, imputation of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
@@
How much stratified sampling, PPI, and confusion matrix-based methods reduce uncertainty relative to classical estimates depends on the quality of the map product. In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum. In PPI (and the Olofsson method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low. In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by $1.2$.
Stratification vs. PPI vs. Olofsson.
Comparing stratified sampling, PPI, and the Olofsson method, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product. Recall Table 2, which summarizes the conditions under which each method is most appropriate. Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI. This is a result of well-designed strata based on a remote sensing map of forest disturbance categories. PPI and the Olofsson method yielded comparable interval widths to each other. This is expected, as the Olofsson method yields similar confidence intervals as PPI when $N$ is large (Appendix A).
@@
An advantage of PPI compared to stratification, then, is its flexibility. While stratification requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset. Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain valid confidence intervals may result in the best performance across a variety of downstream use cases. Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
@@
Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty). We recommend that producers of machine learning-generated remote sensing maps use a holdout ground truth dataset for uncertainty quantification. The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate. Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences. In the absence of a provided holdout set (currently, releasing ground truth data is not the norm), we recommend that map users generate their own ground truth data for the variable of interest if possible.
@@
All of the estimation methods described in this paper (except for naive imputation) are limited to areas in which ground truth data is available. However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias. Furthermore, we can only ensure statistical validity of our estimates when the ground truth points are randomly sampled. Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys. Another limitation is that our estimates only quantify uncertainty in the target variable map product, and do not account for noise in either the ground truth data or (in the regression use cases) the covariates. We leave these investigations to future work.

### v3

@@ 2.4 Post-stratified area estimator
$SE_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
@@
$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot SE_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot SE_{j,\text{post}}).$
Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$. Thus, we can also perform area estimation using methods such as PPI and stratified estimation. If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval is identical to the stratified confidence interval.
@@ 3 Area estimation
In this section, we demonstrate uncertainty quantification for two area estimation use cases. First, we simulate estimating maize area in Iowa in 2022, artificially degrading the map product quality relative to simulated “ground truth” points, in order to observe the effect of map product noise and bias on the area estimate under different methods. We will see through this simulation the importance of bias correction before using maps to estimate total area. We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015. We compare the classical, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.
@@ 3.1 Simulation: Maize area in Iowa
In each trial, we sample $n=100$ “ground truth” points uniformly at random from the $N=100000$ map product points. Initially, we set the map product and ground truth equal to each other at the 100 ground truth locations. We then experiment with degrading the map product quality as follows. We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), classical, and PPI estimators. We average results over 100 trials.
@@ Results
As we add increasing amounts of random noise to the map, all confidence intervals for maize area remain centered at the true maize fraction of 0.35 (Figure 5). This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same. The classical confidence interval stays constant because it only uses ground truth points. The map-only confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product. However, the map-only confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
@@
When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals. This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error. As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the classical confidence interval at noise probability $p=1$. (At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth samples.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$). Thus, the effectiveness of PPI is dependent on the quality of the map product. The post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI.
@@
Figure 5: Simulation estimating maize fraction under different levels of map noise using $N=100000$ map product points and $n=100$ “ground truth” points. Post-stratified and PPI results are overlapping. Left: Maize fraction 95% confidence intervals at different noise levels. Right: Maize fraction effective sample size at different noise levels. (At noise level $p$, about $p$ fraction of the points are resampled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 (the true maize fraction) and non-maize with probability 0.65.)
@@
Figure 6: Simulation estimating maize fraction 95% confidence intervals under different levels of map bias using $N=100000$ map product points and $n=100$ “ground truth” points. Post-stratified and PPI confidence intervals are overlapping. (Noise is fixed at 0.4 so about 40% of the points are resampled. At bias $p$, each resampled point is labeled as maize with probability $p$ and non-maize with probability $1-p$.)
@@
The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 6. Again, the classical confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points. As bias increases, the fraction of map product points labeled as “maize” increases. As a result, the map-only estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35. The map-only confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product. In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product. The PPI confidence intervals are smaller than the classical intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level). Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points. Again, the post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
@@
We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7). As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and post-stratified confidence intervals. Thus, the left panel of Figure 7 resembles an inverted version of the left panel of Figure 5.
@@
Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points. Post-stratified and PPI confidence intervals are overlapping.
@@
As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly. (To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as model accuracy increases. In the right panel of Figure 7, a model accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35. This shows that even a map with high model accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
@@
Figure 9: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points. The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015. The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9). The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples. The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$. (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth. In particular, the stratified estimator is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
In this case, the map-only (pixel-counting) estimate — using only the map to estimate deforestation without any bias correction via ground truth labels — falls within the other confidence intervals, suggesting that the NASA Global Forest Cover Change map has small bias for estimating deforestation area. However, we emphasize that this would not have been clear without the ground truth data, and ground truth data is still necessary to check the reliability of the map product for deforestation analysis.
@@ 4 Estimating regression coefficients
In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021. We compare the classical method, map-only estimator, PPI, stratified estimator, and stratified PPI. Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
@@ Results
The general degree of uncertainty in estimating the coefficients from our non-map-only methods are as follows. The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope. Thus, deforestation is negatively associated with all covariates except for elevation. The elevation coefficient confidence intervals intersect with zero for all non-map-only methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate. The most significant improvement is for the slope covariate. Unlike in our previous area estimation examples, the stratified estimator (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers. Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads. The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
For slope and distance from rivers, the map-only estimated coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points. The map-only coefficient for slope is about 0, and the map-only coefficient for distance from rivers is about $-0.010$. In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero. However, if we had used only the remote sensing map, we would not have discovered the negative relationship between slope and deforestation. Thus, using only the map product could result in misleading conclusions due to biases in the deforestation map product that are correlated with covariates (e.g., the map product overestimates deforestation at locations with high slope, or underestimates deforestation at locations with low slope).
Figure 11: Brazilian Amazon deforestation logistic regression coefficient 95% confidence intervals using $N=963548$ map product points and $n=1386$ ground truth points. The PPI methods result in the smallest confidence intervals, while stratification is not generally effective. For slope and distance from rivers, the map-only coefficient falls outside the confidence intervals for all methods that use ground truth points.
@@
Figure 13: US tree cover linear regression coefficient 95% confidence intervals using $N=983238$ map product points and $n=983$ ground truth points. PPI outperforms the classical method. For elevation, the map-only coefficient falls outside the PPI confidence interval.
@@
For elevation, the map-only coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval. This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations). Similar to the deforestation logistic regression use case, using only the map product could result in misleading conclusions. If we had used only the remotely sensed tree cover predictions, we would have overestimated the magnitude of the relationship between elevation and tree cover due to map errors.
@@ 5 Discussion
Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures. In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products. We find that stratified, post-stratified, and prediction-powered inference estimates result in lower uncertainty than classical estimates using only ground truth data sampled. Unlike map-only estimates using the map product alone, these estimates are guaranteed to be unbiased.
Bias is the main issue for map-only estimates.
Our experiments illustrated that a map-only estimate of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias. In our simulation estimating maize fraction, map-only estimates deviated from the true maize fraction of 0.35 as the bias of the crop type map increased. Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
In our experiments on real data, the map-only estimate for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product. However, the map-only estimates of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods. The map-only estimator would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does. Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients. Similarly, the map-only estimate of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
@@
How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the classical estimator depends on the quality of the map product. In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum. In PPI (and the post-stratified method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low. In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by 1.2×1.2$\times$1.2 ×.
Stratification vs. Post-stratification vs. PPI.
Comparing stratified, post-stratified, and PPI estimators, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product. Recall Table 2, which summarizes the conditions under which each method is most appropriate. Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI. This is a result of well-designed strata based on a remote sensing map of forest disturbance categories. PPI and the post-stratified estimator yielded comparable interval widths to each other. This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix A).
@@
An advantage of PPI compared to the stratified estimator, then, is its flexibility. While the stratified estimator requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset. Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain confidence intervals may result in the best performance across a variety of downstream use cases. Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
@@
Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty). We recommend that producers of machine learning-generated remote sensing maps use a (simple or stratified) randomly sampled holdout ground truth dataset for uncertainty quantification. The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate, provided that it is randomly sampled. Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences. In the absence of a provided holdout set (currently, releasing ground truth data is not the norm), we recommend that map users generate their own ground truth data for the variable of interest if possible.
@@
All of the estimation methods described in this paper (except for map-only estimators) are limited to areas in which ground truth data is available. However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias. Furthermore, we can only ensure that our estimates are unbiased when the ground truth points are randomly sampled. Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys. Another limitation is that our estimates only quantify uncertainty in the target variable map product, and do not account for noise in either the ground truth data or (in the regression use cases) the covariates. We leave these investigations to future work.